OpenStudy (anonymous):

In the chapter of Applications of Quadratic Equations... A girl's age in 12 years' time is the square of her present age. Find her present age. Help Plz!!!!

OpenStudy (anonymous):

Let the girls age be x. x + 12 = x^2 -x^2 + 12 + x = 0 Solve using quadratic formula now.

OpenStudy (anonymous):

 i subtracted 12 because the question asked the present not her age in 12 years, but we could use it as given. . USE QUADRATIC FORMULA to find the value of x. .